# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v2

@@ D.5 Comparison against TAM
GFACS ($T=4$) 15.77 1.07 24.95 3.23
GFACS ($T=10$) 15.72 1.40 24.87 5.21
We compare GFACS and DeepACO against CVRP instances used in TAM [24]. These instances have a different capacity constraint from those we used for the main experiment (Table 1). Please refer to [24] for more details regarding the instance generation. Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.
D.6 Ablation study
Table 10: Results of the ablation study. The mean and standard deviation are reported based on three independent runs. The reported gap is calculated against Concorde [2] for TSP and PyVRP [77] for CVRP.
@@ D.6 Ablation study
    + VarGrad [82] 16.8570 $\pm$ 0.0086 1.88 64.2771 $\pm$ 0.0293 2.09 51.9504 $\pm$ 0.1222
In this section, we present the comprehensive results of our ablation study, incorporating standard deviations from multiple independent runs (three runs) to ensure statistical significance. The findings demonstrate that each component of our technical proposals contributes to improvement, supported by empirical and statistically rigorous evidence.
@@ Appendix E Symmetric solutions of combinatorial optimization
Appendix F Extended related works
@@ F.1 Deep constructive policy for neural combinatorial optimization
Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution. This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing [57, 41], graph combinatorial optimization [30, 1, 84], and scheduling [81].
In deep learning, various techniques adopt this constructive policy framework. Notably, the attention model (AM) [41] has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training. Subsequent advancements in AM have included enhanced decoder architecture [67, 78, 47], improved REINFORCE [76] baseline assessments [43, 37], leveraging symmetries for sample efficiency [32], the application of population-based reinforcement learning (RL) [21], ensemble-based learning [27, 86] and refined solution-building methods [16]. Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally [36, 24, 80]. Additionally, AM-driven constructive policies have been effectively applied in hardware routing [34, 35] and placement tasks [60, 31].
Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains. Starting with the influential works by Khalil et al. [30], which employed deep Q networks [55] for graph policy training, creating solutions constructively, Ahn et al. [1] introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems. Zhang et al. [84] further developed this concept, training their MDP with forward-looking GFlowNets [59].
Our research aligns with the deep constructive policy framework, where GFlowNets constructively build actions from an initial empty solution ($s_{0}$) to a final complete solution ($s_{n}$). Additionally, our methods include a heatmap-based approach, detailed in Section F.3.
@@ F.2 Deep improvement policy for neural combinatorial optimization
The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt [11]. Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations. Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning [12, 22, 48, 49, 50, 33]. Also, recent works have explored active search to refine solutions via policy adjustments [23, 66] or sampling the conditions from a continuous latent space [7]. Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes [69]. This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods. Additionally, our method, which integrates local search into the training loop, could see further enhancements if advancements in this area are incorporated into our training procedures.
@@ F.3 Heatmap-based neural combinatorial optimization
The heatmap-based method [29] entails the acquisition of a graph representation, referred to as the “Heatmap," from input graph instances. This Heatmap serves as a prior distribution, influencing the generation of a solution. However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search [29], dynamic programming [42], Monte Carlo tree search [20, 62, 70] or greedy search with a guide using an objective function [44]. The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies. It operates at an abstract level, free from the constraints inherent in generating actual combinatorial solutions.
@@
Broader Impact
We proposed a new probabilistic approach that can be trained using off-policy learning and applied to various combinatorial optimization problems, such as vehicle routing and scheduling. This method is relevant to industrial engineering, particularly in optimizing transportation and manufacturing processes, and can potentially extend to scientific discovery through chemical structure and biological sequence optimization. It is important to consider the ethical implications of probabilistic optimization methods, as they carry the potential for misuse in fields such as harmful drug discovery. While these algorithms offer several benefits, their application should be carefully monitored to ensure responsible and ethical use, minimizing the risk of unintended negative consequences.

### v4

@@ D.4 Comparison against TAM
GFACS ($T=4$) 15.77 2.91 24.95 3.23
GFACS ($T=10$) 15.72 3.88 24.87 5.21
We compare GFACS and DeepACO against CVRP instances used in Two-stage Divide Method (TAM; Hou et al.,, 2023). These instances have a different capacity constraint from those we used for the main experiment (Table 2). Please refer to Hou et al., (2023) for more details regarding the instance generation. Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (the number of ants for ACO) to 16.
D.5 Full results of the ablation study
Table 10: Results of the ablation study. The mean and standard deviation are reported based on three independent runs. The reported gap is calculated against Concorde (Applegate et al.,, 2006) for TSP and PyVRP (Wouda et al.,, 2024) for CVRP.
@@ D.5 Full results of the ablation study
    + VarGrad Zhang et al., 2022b 16.8570 $\pm$ 0.0086 1.88 64.2771 $\pm$ 0.0293 2.09 51.9504 $\pm$ 0.1222
In this section, we present the comprehensive results of our ablation study, including standard deviations from multiple independent runs (three runs) to ensure statistical significance. Table 10 demonstrates that each component of our technical proposals contributes to improvement, supported by empirical and statistically rigorous evidence.
D.6 Results without local search on TSP and CVRP
Table 11: Results obtained with and without local search (LS) at test time on TSP and CVRP are reported. The values represent the objective value averaged over 128 test instances and further averaged across three independent runs.
Method TSP200 TSP500 TSP500 CVRP500
ACO w/o LS 14.19 24.05 42.95 95.24
DeepACO w/o LS 11.59 18.83 32.12 73.20
GFACS w/o LS 12.63 21.72 34.04 81.03
ACO w/ LS 10.91 17.38 28.45 64.62
DeepACO w/ LS 10.77 16.84 28.44 64.49
GFACS w/ LS 10.75 16.80 28.32 64.19
We conduct an additional ablation study to assess the impact of local search at test time. Note that we use the same policies as in Section 5.1 or Section 5.2, which were trained with local search. The results are presented in Table 11. As expected, DeepACO (RL-based) performs better than ours when local search is not applied. This outcome can be attributed to two factors: (1) The RL-based method tends to generate non-diverse low-energy samples, whereas the GFlowNet-based method (GFACS) produces a more diverse set of candidates; and (2) Due to the energy reshaping (Eq. (7) in Section 4.1), the policy is optimized with respect to a modified energy map that assumes the application of local search, resulting in suboptimal performance when local search is omitted. This second factor also explains why GFACS outperforms other methods when local search is applied.
D.7 Runtime analysis
Table 12: Average running time (in seconds) to solve a single instance for Table 1.
$N$ ACO DeepACO GFACS
TSPLib 100-299 1.3 1.3 1.3
300-699 11.2 10.5 10.8
700-1499 89.3 72.5 77.0
CVRPLib 100-299 15.5 15.5 15.5
300-699 34.2 35.1 35.5
700-1001 60.0 59.8 61.8
For completeness, in Table 12, we report the runtime for each algorithm on the real-world instances (Table 1 in Section 5.1). Although all algorithms have similar inference times, DeepACO and GFACS occasionally run faster than ACO. This is attributed to the increased complexity of real-world instances: both DeepACO and GFACS benefit from an NN-based amortized prior that enables them to start closer to a local optimum, resulting in earlier termination of the local search. This also explains why DeepACO is slightly faster than GFACS, as it produces a more locally optimized initial solution.
@@ Appendix E Symmetric solutions of combinatorial optimization
Appendix F Related works
@@ F.1 Deep constructive policy for neural combinatorial optimization
Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution. This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing (Nazari et al.,, 2018; Kool et al.,, 2019), graph combinatorial optimization (Khalil et al.,, 2017; Ahn et al.,, 2020; Zhang et al.,, 2023), and scheduling (Zhang et al.,, 2020).
In deep learning, various techniques adopt this constructive policy framework. Notably, the attention model (AM; Kool et al.,, 2019) has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training. Subsequent advancements in AM have included enhanced decoder architecture and decoding procedure (Son et al.,, 2024; Xin et al.,, 2021; Luo et al.,, 2023; Drakulic et al.,, 2023), improved REINFORCE (Williams,, 1992) baseline assessments (Kwon et al.,, 2020; Kim et al.,, 2022), leveraging symmetries for sample efficiency (Kim et al., 2024a, ), the application of population-based reinforcement learning (RL) (Grinsztajn et al.,, 2023), ensemble-based learning (JIANG et al.,, 2023; Zhou et al.,, 2024). Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally (Kim et al.,, 2021; Hou et al.,, 2023; Ye et al.,, 2024; Zheng et al.,, 2024).
Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains. Starting with the influential works by Khalil et al., (2017), which employed deep Q networks (Mnih et al.,, 2013) for graph policy training, creating solutions constructively, Ahn et al., (2020) introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems. Zhang et al., (2023) further developed this concept, training their MDP with forward-looking GFlowNets (Pan et al.,, 2023).
Our research aligns with the deep constructive policy framework, where we build a solution from an initial empty solution ($s_{0}$) based on the policy trained with GFlowNets. Additionally, our method is included in the heatmap-based approaches, detailed in Section F.3.
@@ F.2 Deep improvement policy for neural combinatorial optimization
The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt (Croes,, 1958). Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations. Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning (d O Costa et al.,, 2020; Hottung and Tierney,, 2020; Ma et al.,, 2021, 2022, 2023; Kim et al.,, 2023). Also, recent works have explored active search to refine solutions via policy adjustments (Hottung et al.,, 2022; Son et al.,, 2023) or sampling the conditions from a continuous latent space (Chalumeau et al.,, 2023). Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes (Sun et al.,, 2023). This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods. Additionally, our method, which integrates local search into the training loop, could see further enhancements if advancements in this area are incorporated into our training procedures.
@@ F.3 Heatmap-based neural combinatorial optimization
The heatmap-based method (Joshi et al.,, 2021) entails the acquisition of a graph representation, referred to as the “Heatmap," from input graph instances. This Heatmap serves as a prior distribution, influencing the generation of a solution. However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search (Joshi et al.,, 2021), dynamic programming (Kool et al.,, 2022), Monte Carlo tree search (Fu et al.,, 2021; Qiu et al.,, 2022; Sun and Yang,, 2023; Xia et al.,, 2024) or greedy search with a guide using an objective function (Li et al.,, 2023). The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies. It operates at an abstract level, free from the constraints inherent in generating actual combinatorial solutions.
@@
F.4 GFlowNets for neural combinatorial optimization
Originally proposed as a framework for learning policies over discrete compositional variables (Bengio et al., 2021a, ), GFlowNets have since been applied to combinatorial optimization problems (Zhang et al.,, 2023; Zhang et al., 2022b, ). They offer unique advantages over reinforcement learning methods, particularly in generating diverse candidate solutions and enabling off-policy training. Although prior studies have explored the diversity aspect (Zhang et al.,, 2023; Zhang et al., 2022b, ), our work is, to our knowledge, the first to empirically validate the benefits of off-policy training in combinatorial optimization settings.
To underscore our contribution, we conceptually compare our method against two recent approaches that incorporate off-policy training for GFlowNets: local search GFlowNets (LS-GFN) (LS-GFN; Kim et al., 2024b, ) and the hindsight-like off-policy strategy proposed by Zhu et al., (2023). Note that, in Section 5.4, we provide an empirical comparison of our approach with FL-DB (Zhang et al.,, 2023) and VarGrad (Zhang et al., 2022b, ).
The key difference between LS-GFN and our approach is that our method addresses conditional inference problems. This requires training conditional GFlowNets that operate under distinct reward landscapes for different conditions, whereas LS-GFN is designed for unconditional problems. Moreover, the local search operators themselves differ. LS-GFN uses a back-and-forth mechanism that partially backtracks trajectories and reconstructs them using its own policy, thereby limiting exploration to sequences sharing identical root segments. In contrast, our local search methods, such as the 2-opt operator for the TSP, explore the solution space within a Hamming ball, allowing for more diverse exploration without the constraint of shared initial sequences. Furthermore, our destroy-and-repair local search method (Appendix C) can be viewed as an extension of LS-GFN, incorporating symmetricity-aware destruction to mitigate the limitations of the back-and-forth operator.
In Zhu et al., (2023), they tries to model a family of policies using hypernetworks to tackle multi-objective optimization problems. They also propose a hindsight-like off-policy strategy, which enables sharing of off-policy samples among policies configured with varying objective mixing coefficients, thereby enhancing sample efficiency during training. In contrast, our method leverages hindsight observations from a (deterministic) local search (Section 4.1). Specifically, we adjust the energy of an original sample by incorporating the energy after local search, thereby quantifying how readily the sample can be converted into a low-energy solution.